A copper wire has an original length of 1.0m and a diameter of 1.1mm. What is the cross-sectional area of the wire?


1
Expert's answer
2020-11-17T13:03:25-0500

The cross-sectional area of the wire is


"A=\pi\frac{d^2}{4}\\A=\pi\frac{1.1^2}{4}=0.95\ mm^2"